Find the point at which the line intersects the given plane. x = y – 2 = 4z; 4x – y + 2z = 12 (x, y, z) = ( (x, y, z) = D
Answer #1

given x = Y-2 = 42 plane 4x – 4+22.512 - put o in 464-2) –4 +2z = 12 44-8-Y + 22 =12 34+22 = 20 ~ from 0¢ ® Y-42= 2 = 34–422=